OpenStudy (anonymous):

a carpenter uses 26ft of lumber to build 2 frames. If he is going to build 5 door frames, how much lumber does he need?

OpenStudy (anonymous):

2 FRAMES NEED 26FT 5 FRAMES NEED 26/2 * 5= 65

OpenStudy (anonymous):

\[26/2 \times 5 = 65\] correct?!